# Attention Projection Mixing with Exogenous Anchors

### v2

@@ Abstract
Cross-layer reuse of early attention projections can improve optimization and data efficiency, but it creates a structural conflict: the first layer must simultaneously act as a stable, reusable anchor for all deeper layers and as an effective computational block. We show this ”first-layer tension” is a hidden limiter of internal-anchor designs. We propose ExoFormer, which resolves the conflict by learning exogenous anchor projections outside the sequential layer stack, decoupling the anchor role from computational refinement. We introduce a unified normalized mixing framework that mixes queries, keys, values, and gate logits using learnable coefficients (exploring coefficient granularities: elementwise/headwise/scalar), and we show that normalizing anchor sources is key to stable reuse. ExoFormer variants consistently outperform their internal-anchor counterparts, and the dynamic variant yields $\sim\!1.5$ downstream accuracy points while matching validation loss using $\sim\!1.5\times$ fewer tokens than Gated Attention. We explain this efficacy via an Offloading Hypothesis: external anchors preserve essential token identity, allowing layers to specialize exclusively in refinement. We release code and models to facilitate future research.
@@ 1 Introduction
However, existing solutions largely operate in isolation and leave key questions unanswered. ResFormer (Zhou et al., 2025) focuses solely on residualizing values, leaving the potential reuse of other attention components—queries, keys, and gating logits—unexplored.
@@
(a) Validation loss for $\sim\!450M$ parameter models
(b) Validation loss for $\sim\!1B$ parameter models
Analyzing architectures that reuse first-layer projections reveals a fundamental tension. The first layer is forced to serve two roles: (1) as a stable, reusable anchor for all deeper layers and (2) as an effective computational block for progressive feature transformation. This forced coexistence inherently limits effectiveness in both roles.
Our primary contribution is ExoFormer, which resolves this tension by learning a set of dedicated exogenous anchor projections outside the sequential layer stack. The internal-anchor version, which we term NuResFormer (Normalized unified), serves as a comparative foundation. Decoupling of the two roles proves consistently beneficial: every ExoFormer variant outperforms its corresponding NuResFormer counterpart in perplexity while remaining competitive in downstream accuracy despite no increase in width or depth, with the dynamic variant achieving the best results. We explain this high performance via an Offloading Hypothesis: the external anchor assumes the role of preserving token identity, allowing the sequential layers to specialize almost exclusively in refinement.
@@ 2 Related Work
Most closely related to our work—though orthogonal in direction—is MUDDFormer (Xiao et al., 2025), which proposes Multiway Dynamic Dense (MUDD) connections. MUDDFormer decouples the input to each Transformer block into four streams (query, key, value, residual) and dynamically aggregates outputs from all preceding layers using context‑specific weights generated by a small MLP.
While techniques mentioned above and our work focus on enhancing residual pathways, they operate at different levels and can coexist. MUDDFormer primarily addresses the input to the attention projections. In contrast, our unified mixing framework focuses on mixing the attention projections themselves. Our approach can be seen as a form of residual blending after projection, whereas MUDDFormer enriches the input before projection.
@@ 3.1 Preliminaries and Notation
We consider three coefficient granularities for $\lambda$: Scalar (S) (a single coefficient shared across all channels), Headwise (H) (one coefficient per head, broadcast across $d_{k}$), and Elementwise (E) (one coefficient per channel).
@@ 3.2 Multi-Head Attention
We describe attention as a sequence of stages, with two standard enhancements integrated implicitly: rotary position embeddings (RoPE) (Su et al., 2023) and query/key normalization (QKNorm) (Henry et al., 2020). For simplicity, the multi-head mechanism is presented using unified tensors with an implicit head dimension.
@@ 3.3 Gated Attention
$G_{n}=H_{n-1}W^{G}_{n},\qquad W^{G}_{n}\in\mathbb{R}^{d_{\text{model}}\times(hd_{k})}.$ (5)
@@ 3.5 First Layer Tension and ExoFormer
While these roles appear misaligned—universal anchors favor invariance, while progressive computation necessitates change—they can theoretically coexist because it is an optional pathway modulated by learned mixing coefficients ($\lambda_{n,1},\lambda_{n,2}$). However, this forced coexistence places pressure the first layer to make compromises, inherently limiting its effectiveness in both roles. This tension is empirically supported by the analysis in Figure 7, showing that NuResFormer’s first layer adopts a permissive gating policy compared to standard baselines, indicating a compromise on selectivity in favor of serving as a stable anchor.
Our most successful approach, termed ExoFormer (where “Exo” signifies an external, outside-the-system anchor), instantiates the general framework of Section 3.4 with the dedicated exogenous projections of Eq. (9).
This decoupling proves beneficial: in our experiments, every ExoFormer variant outperformed its corresponding NuResFormer counterpart in perplexity, despite no change to depth or width.
@@ 4.1 Experimental Setup
All models use a modern pre-normalized Transformer architecture with SwiGLU activations (Shazeer, 2020), QKNorm (Henry et al., 2020), and rotary position embeddings (Su et al., 2023). We follow prior work by initializing projection and classification layers to zero (Yang et al., 2022), removing bias terms (except for the dynamic mixing module) (Chowdhery et al., 2022), applying z-loss regularization (debrébisson2016zlossshiftscaleinvariant), and disabling dropout. Training is performed on the FineWeb-Edu dataset (Penedo et al., 2024).
Table 1: Performance comparison of model variants on 6 multiple-choice downstream tasks. The parameter counts are: 453M for models with gated attention, 454M for those without it, and 457M for models that incorporate an external anchor. Naïve Combination refers to the unmodified addition of Gated Attention and ResFormer. Dynamic ExoFormer achieves the highest overall accuracy and the lowest perplexity, demonstrating the advantages of decoupling the anchor from the first layer and using context-aware mixing. All models use full normalization on anchors unless noted otherwise. The notation V, G denotes ablations where only value and gate projections are mixed; queries and keys are computed solely from the current layer. All model labels refer specifically to Anchor Normalization (e.g., “Only Q/K Norm” means RMSNorm is applied only to anchor Q/K sources).
@@
We optimize matrix parameters with Muon (specifically the Polar Express variant; Amsel et al. (2025)), applying a cautious weight decay of 0.1 (Chen et al., 2025), while 1D parameters are trained using AdamW without weight decay. Muon was utilized because it represents a state-of-the-art optimizer for large-scale LLM training (Liu et al., 2025). Gradient norms are clipped to 1.0, and all models follow the same optimization setup, to ensure fair comparison across architectures. Training uses a global batch size of 262,144 tokens and a sequence length of 2,048.
We apply linear learning-rate warmup for the first 1,000 steps, with peak learning rates of $3\times 10^{-3}$ for AdamW and $1\times 10^{-2}$ for Muon. Following warmup, learning rates decay linearly to 10% of their peak values for 20% of the total training steps (hägele2024scalinglawscomputeoptimaltraining). Full hyperparameters are provided in the Appendix. All experiments are conducted on a single NVIDIA H100 80GB GPU using native BF16 precision, with FlashAttention (Dao et al., 2022) enabled.
@@
We report results on 6 multiple-choice benchmarks: arc_challenge, arc_easy (Clark et al., 2018), hellaswag (Zellers et al., 2019), openbookqa (Mihaylov et al., 2018), piqa (Bisk et al., 2019), and winogrande (Sakaguchi et al., 2019).
@@ 4.2 The Role of Anchor Normalization
We hypothesize that the naïve combination underperforms for the above reason. When applied to unbalanced mixtures, $\sigma(G)$ must compensate for distributional mismatch rather than execute content-aware filtering, resulting in instability during training (see Figure 2(a)) and worse downstream accuracy than ResFormer in isolation.
@@ 4.3 Extending Mixing to Q, K, and G Pathways
Consistent with prior observations (Zhou et al., 2025), we find a clear hierarchy of stability. Models attempting to use unnormalized Q/K residuals without QKNorm proved impossible to optimize, exhibiting divergent loss in preliminary runs. Critically, adding QKNorm alone stabilizes training and recovers baseline performance. We hypothesize this is because QKNorm compresses the scale of $Q$ and $K$, mitigating the distributional mismatch that arises when injecting early-layer routing signals into deeper layers. Furthermore, adding explicit RMSNorm to the Q/K anchor sources on top of QKNorm yields the best results, enabling positive reuse.
@@ 4.4 Mix-Compress-Refine Theory, Gated Attention, and ResFormer
Similarly, in models employing residual mixing, the residual pathways may provide an implicit, learned alternative to harsh sink-based filtering. The learned blending could reduce the model’s reliance on extreme, attention-sink-driven compression to isolate useful signals. This is supported by the milder compression valley observed in such models (Figure 4(c)) and their notable reduction in attention sink magnitude (Table 2).
@@ 4.5 The Offloading Hypothesis: Specialization via Exogenous Anchors
The architectural decoupling in ExoFormer enables an interesting functional specialization, which we formalize as the Offloading Hypothesis. By providing a dedicated, high-fidelity source of token identity, the exogenous anchor allows the sequential layers to offload the burden of preserving static features and specialize almost exclusively in the final ”refinement” stage (Stage 3) of the Mix-Compress-Refine cycle.
Evidence for this specialization is clearly visible in the token-similarity trajectory (Figure 4(b)). All models exhibit a characteristic pattern: similarity rises to an early maximum (Stage 1: mixing), dips to a local minimum (Stage 2: compression), rises again, and finally falls in the last layer (Stage 3: refinement). We interpret the local minimum as the onset of Stage 3—the refinement phase dedicated to preparing discriminative representations for next-token prediction. Crucially, ExoFormer variants and NuResFormer spend approximately two-thirds of their layers in this specialized refinement stage, the highest proportion of any model, while the baseline spends only one-third.
This specialization is possible because exogenous anchors fundamentally reconfigure information management within the Transformer. In a standard architecture, the residual stream must concurrently carry two conflicting types of information: (1) distinct, static features that preserve token identity (”What token am I?”), and (2) transformed, task-relevant features that evolve through the layers to support next-token prediction. Over-smoothing in standard models catastrophically loses the first type, crippling the model’s ability to route information effectively. ExoFormer circumvents this tension by introducing a permanent reference library—the exogenous anchor—that is reinjected at every layer, guaranteeing access to high-fidelity token identity.
@@
    The Anchor Specializes in ”Identity Preservation.” By providing a guaranteed, high-fidelity source of token distinctiveness at every layer, the exogenous anchor relieves the main residual stream from the duty of preserving static token features. The dedicated anchor parameters ($W^{Q}_{\text{anc}},W^{K}_{\text{anc}},W^{V}_{\text{anc}},W^{G}_{\text{anc}}$) can specialize exclusively in maximizing variance and distinctiveness, optimized to serve as a universal reference.
@@
    The Sequential Layers Specialize in ”Progressive Computation.” With identity information externally supplied, the internal layers are liberated from the constraint of preserving early-layer geometry. They are no longer required to maintain a stable, high-fidelity record of token identity within the residual stream—a task that inherently conflicts with the need to transform and abstract features.
@@ 5 Conclusion
While our contribution is architectural—focused on efficiency and structure rather than specific content generation—it facilitates the deployment of more capable AI systems that could be employed maliciously. Additionally, while our ablations suggest robustness, there is inherent uncertainty in how the ”offloading” mechanism will scale to significantly larger parameter counts (e.g., 100B+), which may introduce novel failure modes or safety challenges not observed in the 450M–1B range studied here.
@@ C.1 Parameter Overhead
For values ($L=32,d=1024,d_{\text{DM}}=16$), this corresponds to approximately 1.2% overhead for static ExoFormer and 1.3% for dynamic ExoFormer—a modest increase given the observed performance gains.
@@ C.3 Efficiency-Performance Trade-off
The FLOPs analysis presented above provides a lower-bound estimate of the computational overhead. In practice, the real-world slowdown may be higher (we observe an increase in latency per token of approximately 8–15% for our reference implementation prioritizing modularity and interpretability over kernel fusion, compared to the theoretical  1.3% FLOPs increase).

### v3

@@ Abstract
Cross-layer reuse of early attention projections can improve optimization and data efficiency, but it creates a structural conflict: the first layer must simultaneously act as a stable, reusable anchor for all deeper layers and as an effective computational block. We demonstrate that this tension constrains the performance of internal-anchor designs. We propose ExoFormer, which resolves the conflict by learning exogenous anchor projections outside the sequential layer stack. We introduce a unified normalized mixing framework that mixes queries, keys, values, and gate logits using learnable coefficients (exploring coefficient granularities: elementwise, headwise, and scalar), and we show that normalizing anchor sources is key to stable reuse. ExoFormer variants consistently outperform their internal-anchor counterparts, and the dynamic variant yields $\sim\!1.5$ downstream accuracy points while matching validation loss using $\sim\!1.5\times$ fewer tokens than Gated Attention. We explain this efficacy via an Offloading Hypothesis: external anchors preserve essential token identity, allowing layers to specialize exclusively in feature transformation. We release code and models to facilitate future research.
@@ 1 Introduction
However, existing solutions largely operate in isolation and leave key questions unanswered. ResFormer (Zhou et al., 2025) focuses solely on residualizing values, leaving the potential reuse of other attention components (queries, keys, and gating logits) unexplored.
@@
Figure 2: (Left) Validation loss for $\sim\!450M$ parameter models. (Right) Validation loss for $\sim\!1B$ parameter models.
Analyzing architectures that reuse first-layer projections reveals a fundamental tension. The first layer is forced to serve two roles: (1) as a stable, reusable anchor for all deeper layers and (2) as an effective computational block for progressive feature transformation. This dual objective inherently limits effectiveness in both roles.
Our primary contribution is ExoFormer, which resolves this tension by learning a set of dedicated exogenous anchor projections outside the sequential layer stack. The internal-anchor version, which we term NuResFormer (Normalized unified), serves as a comparative foundation. Decoupling of the two roles proves consistently beneficial: every ExoFormer variant outperforms its corresponding NuResFormer counterpart in perplexity while remaining competitive in downstream accuracy despite no increase in width or depth. We explain this high performance via an Offloading Hypothesis: the external anchor assumes the role of preserving token identity, allowing the sequential layers to specialize almost exclusively in refinement.
@@ 2 Related Work
Most closely related to our work, though orthogonal in direction, is MUDDFormer (Xiao et al., 2025), which proposes Multiway Dynamic Dense (MUDD) connections. MUDDFormer decouples the input to each Transformer block into four streams (query, key, value, residual) and dynamically aggregates outputs from all preceding layers using context‑specific weights generated by a small MLP.
While the techniques mentioned above and our work focus on enhancing residual pathways, they operate at different levels and can coexist. MUDDFormer primarily addresses the input to the attention projections. In contrast, our unified mixing framework focuses on mixing the attention projections themselves. Our approach can be seen as a form of blending after projection, whereas MUDDFormer enriches the input before projection.
@@ 3.1 Preliminaries and Notation
We consider three coefficient granularities for $\lambda$: Scalar (S) (a single coefficient shared across all channels), Headwise (H) (one coefficient per head, broadcast across $d_{k}$), and Elementwise (E) (one coefficient per channel). Elementwise and headwise granularities grant fine-grained control over reuse mechanisms. We further analyze how they impact training and downstream performance.
@@ 3.2 Multi-Head Attention
We describe attention as a sequence of stages, with two standard enhancements integrated implicitly: rotary position embeddings (RoPE) (Su et al., 2024) and query/key normalization (QKNorm) (Henry et al., 2020). For simplicity, the multi-head mechanism is presented using unified tensors with an implicit head dimension.
@@ 3.3 Gated Attention
$G_{n}=H_{n-1}W^{G}_{n},\qquad W^{G}_{n}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{model}}}.$ (5)
@@ 3.5 First Layer Tension and ExoFormer
While these roles appear misaligned (universal anchors favor invariance, while progressive computation necessitates change), they can theoretically coexist because it is an optional pathway modulated by learned mixing coefficients ($\lambda_{n,1},\lambda_{n,2}$). However, this structural constraint places pressure on the first layer to make compromises, inherently limiting its effectiveness in both roles. This tension is empirically supported by the analysis in Figure 7, showing that NuResFormer’s first layer adopts a permissive gating policy compared to standard baselines, indicating a compromise on selectivity in favor of serving as a stable anchor.
Our most successful approach, termed ExoFormer, instantiates the general framework of Section 3.4 with the dedicated exogenous projections of Eq. (9).
@@ 4.1 Experimental Setup
All models use a modern pre-normalized Transformer architecture with SwiGLU activations (Shazeer, 2020), QKNorm (Henry et al., 2020), and rotary position embeddings (Su et al., 2024). We follow prior work by initializing projection and classification layers to zero (Yang et al., 2022), removing bias terms (except for the dynamic mixing module) (Chowdhery et al., 2023), applying z-loss regularization (debrébisson2016zlossshiftscaleinvariant), and disabling dropout. Training is performed on the FineWeb-Edu dataset (Penedo et al., 2024). Specifically, the $\sim\!450\text{M}$ parameter models are trained on $10\text{B}$ tokens, while the $\sim\!1\text{B}$ parameter models are trained on $20\text{B}$ tokens.
Table 1: Performance comparison of model variants on 6 multiple-choice downstream tasks. The parameter counts are: 453M for models with gated attention, 454M for those without it, and 457M for models that incorporate an external anchor. Naïve Combination refers to the unmodified addition of Gated Attention and ResFormer. Dynamic ExoFormer achieves the highest overall accuracy and the lowest perplexity, demonstrating the advantages of decoupling the anchor from the first layer and using context-aware mixing. All models use full normalization on anchors unless noted otherwise. The notation {V,G} denotes ablations where only value and gate projections are mixed; queries and keys are computed solely from the current layer. Prefixes E, H, and S denote elementwise, headwise, and scalar mixing granularity. “Only Q/K Norms” applies RMSNorm solely to anchor queries/keys, while “No Norm” omits anchor normalization.
@@
We optimize matrix parameters with Muon (Polar Express variant (Amsel et al., 2025)), applying a cautious weight decay of 0.1 (Chen et al., 2025), while 1D parameters are trained using AdamW without weight decay. We selected Muon as a state-of-the-art optimizer for large-scale LLM training (Liu et al., 2025), though we anticipate that the observed improvements should be optimizer-agnostic. Gradient norms are clipped to 1.0, and all models follow the same optimization setup, to ensure fair comparison across architectures. Training uses a global batch size of 262,144 tokens and a sequence length of 2,048.
Full hyperparameters are provided in the Appendix. All experiments are conducted on a single NVIDIA H100 80GB GPU using native BF16 precision, with FlashAttention (Dao et al., 2022) enabled.
@@
We report results on 6 multiple-choice benchmarks: arc_challenge, arc_easy (Clark et al., 2018), hellaswag (Zellers et al., 2019), openbookqa (Mihaylov et al., 2018), piqa (Bisk et al., 2020), and winogrande (Sakaguchi et al., 2020).
@@ 4.2 The Role of Anchor Normalization
We hypothesize that the naïve combination underperforms for the above reason. $\sigma(G)$ must compensate for distributional mismatch rather than filtering, resulting in instability during training (Figure 2) and worse downstream accuracy than ResFormer in isolation.
@@ 4.3 Extending Mixing to Q, K, and G Pathways
Consistent with prior observations (Zhou et al., 2025), we find a clear hierarchy of stability. Models attempting to use unnormalized Q/K residuals without QKNorm proved difficult to optimize, exhibiting divergent loss in preliminary runs. Critically, adding QKNorm alone stabilizes training and recovers baseline performance. We hypothesize this is because QKNorm compresses the scale of $Q$ and $K$, mitigating the distributional mismatch that arises when injecting early-layer routing signals into deeper layers. Furthermore, adding explicit RMSNorm to the Q/K anchor sources on top of QKNorm yields the best results, enabling positive reuse.
@@ 4.4 Mix-Compress-Refine Theory, Gated Attention, and ResFormer
Similarly, in models employing residual mixing, the residual pathways may provide an implicit, learned alternative to abrupt sink-based filtering. The learned blending could reduce the model’s reliance on extreme, attention-sink-driven compression to isolate useful signals. This is supported by the milder compression valley observed in such models (Figure 4(c)) and their notable reduction in attention sink magnitude (Table 2).
@@ 4.5 The Offloading Hypothesis: Specialization via Exogenous Anchors
The architectural decoupling in ExoFormer enables an interesting functional specialization, which we formalize as the Offloading Hypothesis. By providing a dedicated, high-fidelity source of token identity, the exogenous anchor allows the sequential layers to offload the burden of preserving static features and specialize almost exclusively in the final “refinement” stage (Stage 3) of the Mix-Compress-Refine cycle.
Evidence for this specialization is clearly visible in the token-similarity trajectory (Figure 4(b)). All models exhibit a characteristic pattern: similarity rises to an early maximum (Stage 1: mixing), dips to a local minimum (Stage 2: compression), rises again, and finally falls in the last layer (Stage 3: refinement). We interpret the local minimum as the onset of Stage 3. Crucially, ExoFormer variants and NuResFormer spend approximately two-thirds of their layers in this specialized refinement stage, the highest proportion of any model, while the baseline spends only one-third.
This specialization is possible because exogenous anchors fundamentally reconfigure information management within the Transformer. In a standard architecture, the residual stream must concurrently carry two conflicting types of information: (1) distinct, static features that preserve token identity (“What token am I?”), and (2) transformed, task-relevant features that evolve through the layers to support next-token prediction. Over-smoothing in standard models catastrophically loses the first type, crippling the model’s ability to route information effectively. ExoFormer circumvents this tension by introducing a persistent external anchor that is reinjected at every layer, guaranteeing access to high-fidelity token identity.
@@
    The Anchor Specializes in “Identity Preservation.” By providing a guaranteed, high-fidelity source of token distinctiveness at every layer, the exogenous anchor removes the requirement for the main residual stream to preserve static token features. The dedicated anchor parameters ($W^{Q}_{\text{anc}},W^{K}_{\text{anc}},W^{V}_{\text{anc}},W^{G}_{\text{anc}}$) can specialize exclusively in maximizing variance and distinctiveness, optimized to serve as a universal reference.
@@
    The Sequential Layers Specialize in “Progressive Computation.” With identity information externally supplied, the internal layers are decoupled from the constraint of preserving early-layer geometry. They are no longer required to maintain a stable, high-fidelity record of token identity within the residual stream, a task that inherently conflicts with the need to transform and abstract features.
@@ 5 Conclusion
While our contribution is architectural (focused on efficiency and structure rather than specific content generation), it facilitates the deployment of more capable AI systems that could be employed maliciously. Additionally, while our ablations suggest robustness, there is inherent uncertainty in how the “offloading” mechanism will scale to significantly larger parameter counts (e.g., 100B+), which may introduce novel failure modes or safety challenges not observed in the 450M–1B range studied here.
@@ C.1 Parameter Overhead
For values ($L=32,d=1024,d_{\text{DM}}=16$), this corresponds to approximately 1.2% overhead for static ExoFormer and 1.3% for dynamic ExoFormer, a modest increase given the observed performance gains.
@@ C.3 Efficiency-Performance Trade-off
The FLOPs analysis presented above provides a lower-bound estimate of the computational overhead. In practice, the real-world slowdown may be higher (we observe an increase in latency per token of approximately 8–15% for our reference implementation prioritizing modularity and interpretability over speed, compared to the theoretical  1.3% FLOPs increase).